OpenStudy (anonymous):

A concrete block is pulled 7.0m across a frictionless surface by means of a rope. The tension in the rope is 40. N, and the network done on the block is 247 J. What angle does the rope make with the horizontal?

OpenStudy (anonymous):

Let \[\theta\] be angle made with the horizontal axis. Work is done by horizontal force \[(7)(40\cos \theta)=247\]